# Joint Learning of Energy-based Models and their Partition Function

### v1

@@ C.6 Joint convexity
~$\widetilde{\mathcal{L}}_{f}(g_{{\bm{w}}},\tau_{{\bm{v}}})$ $\coloneqq\frac{1}{n}\left(\sum_{i=1}^{n}L^{f}_{g_{{\bm{w}}},\tau_{{\bm{v}}}}({\bm{x}}_{i})-g_{{\bm{w}}}({\bm{x}}_{i},{\bm{y}}_{i})\right),$
@@ C.7 Proof of gradient computation
$=\frac{\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))\nabla_{\bm{w}}g_{\bm{w}}({\bm{x}},{\bm{y}})}{\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))}=\frac{\mathbb{E}_{{\bm{y}}\sim q(\cdot|{\bm{x}})}\left[\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))\nabla_{\bm{w}}g_{\bm{w}}({\bm{x}},{\bm{y}})\right]}{\mathbb{E}_{{\bm{y}}\sim q(\dot{|}{\bm{x}})}\left[\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))\right]}$ (87)

### v3

@@ C.6 Joint convexity
$\widetilde{\mathcal{L}}_{f}(g_{{\bm{w}}},\tau_{{\bm{v}}})$ $\coloneqq\frac{1}{n}\left(\sum_{i=1}^{n}L^{f}_{g_{{\bm{w}}},\tau_{{\bm{v}}}}({\bm{x}}_{i})-g_{{\bm{w}}}({\bm{x}}_{i},{\bm{y}}_{i})\right),$
@@ C.7 Proof of gradient computation
$=\frac{\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))\nabla_{\bm{w}}g_{\bm{w}}({\bm{x}},{\bm{y}})}{\sum_{{\bm{y}}\in\mathcal{Y}}q({\bm{y}}|{\bm{x}})\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))}=\frac{\mathbb{E}_{{\bm{y}}\sim q(\cdot|{\bm{x}})}\left[\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))\nabla_{\bm{w}}g_{\bm{w}}({\bm{x}},{\bm{y}})\right]}{\mathbb{E}_{{\bm{y}}\sim q(\cdot|{\bm{x}})}\left[\exp(g_{\bm{w}}({\bm{x}},{\bm{y}}))\right]}$ (82)